Circular No. 7886

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

η CARINAE

M. F. Corcoran, Universities Space Research Association and Goddard Space Flight Center, reports: “Recent observations of this peculiar luminous star with the Proportional Counter Array on the Rossi X-ray Timing Explorer show that the star has undergone an especially strong x-ray flare in the band 2–10 keV. The flare started on Mar. 12, apparently reached x-ray maximum on Apr. 21, has not yet declined to the pre-flare brightness level. The peak x-ray flux in the band 2–10 keV is 1.8 × 10^{−10} erg s^{−1} cm^{−2}; η Car has not been this bright in x-rays since 1997 May 29 (i.e., ∼ 7 months prior to the x-ray eclipse).”

V838 MONOCEROTIS T. Rauch, University of Tübingen; F. Kerber, Space Telescope Eu-

ropean Coordinating Facility, European Southern Observatory (ESO); F. van Wyk, South African Astronomical Observatory (SAAO); and the ESO 2p2 Team at La Silla report a dramatic change in the spectrum of V838 Mon. Observations obtained from Apr. 16 to 22 with the SAAO 1.9-m and ESO 1.5-m telescopes show that strong bands of TiO have emerged and the spectrum now resembles an M5 giant. The photospheric temperature is estimated to be close to 3000 K — a massive drop compared to the 4000 K reported as late as Mar. 9 (cf. IAUC 7855). The molecular features even grew stronger over this period, indicating that the temperature is still decreasing — interpreted as an expansion of the photosphere. Such an expansion is rather more a feature of a late helium-flash object like V4334 Sgr than is typical for a nova.

COMETS C/2002 H3, C/2002 H4, C/2002 H5, C/2002 H6 (SOHO) Further to IAUC 7882, D. Hammer reports his measurements for four

additional Kreutz-sungrazing comets found on C2 images posted at the SOHO website by X. Leprette (C/2002 H3), X.-M. Zhou (C/2002 H4, C/2002 H5), and M. Meyer (C/2002 H6). The reduced astrometry and orbital elements by B. G. Marsden appear on the MPECs cited below.

Comet  2002 UT  α_{2000}  δ_{2000}  MPEC
C/2002 H3 Apr. 17.660  1^{h}49.^{m}4  + 9^{o}40′  2002-H47
C/2002 H4  18.101  1 50.9  + 9 49  2002-H47
C/2002 H5  19.101  1 51.5  + 9 47  2002-H47
C/2002 H6  26.243  2 21.1  +12 31  2002-H49
2002 April 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green